I have a question that states: find (g of f)(x) for the pair of functions. f(x)=x+2 and g(x)=squareroot of x+5, x is greater than or equal -5. I have no clue how to do this!?!
Answer by Nate(3500)

(g of f)(x) means g[f(x)] meaning g(x + 2)
g(x) = sqrt(x + 5)
g(x + 2) = sqrt(x + 2 + 5)
g(x + 2) = sqrt(x + 7)
g[f(x)] >= -5
sqrt(x + 7) >= -5
x + 7 >= 25
x >= 18
With Rule:
x + 7 >= 0
x >= -7